OpenStudy (anonymous):

Stephanie is student teaching at her local high school. She has to help her instructor write a test. She has to write a radical equation where the solution is extraneous. She also has to write a radical equation where the solution is non-extraneous. Help Stephanie write one radical equation where the solution is extraneous and another equation where the solution is non-extraneous. Using complete sentences, explain each step when solving to justify your examples. @satellite73 @iambatman @ganeshie8 @myininaya @Callisto @wio @Kainui @sammixboo

OpenStudy (anonymous):

do they have to be hard?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

as easy as possible please

OpenStudy (anonymous):

then make it easy take some sort of radical expression, anything, and set it equal to a negative number for example \[\sqrt{x-1}=-2\]

OpenStudy (anonymous):

then what do i do?

OpenStudy (anonymous):

it will have no solution if you want to be trickier you can use \(\sqrt{x-1}+2=0\) same thing

OpenStudy (anonymous):

well i need one that the solution will be extraneous and one that would be non extraneous

OpenStudy (anonymous):

\[\sqrt{x-1}+2=0\] will have an "extraneous" solution

OpenStudy (anonymous):

you can to solve it with \[\sqrt{x-1}+2=0\\ \sqrt{x-1}=-2\] square and get \[x-1=4\] so \[x=5\] but if you check \(x=5\) it will not work

OpenStudy (anonymous):

so its an extraneous solution?

OpenStudy (anonymous):

yes it is extraneous whatever that means

OpenStudy (anonymous):

the reason that did not work is because we had \[\sqrt{x-1}=-2\] a negative number if you make the right hand side a positive number it will work you could even use \[\sqrt{x-1}=2\]or something else more imaginative

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw